During the past year, Ironside sold 15 million square yards (units) of carpeting at an average wholesale price of $7.75 per unit. This year, income per capita is expected to surge from $17,250 to $18,750 as the nation recovers from a steep recession. Without any price change, Ironside’s marketing director expects current-year sales to rise to 25 million units.


A. Calculate the implied income arc elasticity of demand. B. Given the projected rise in income, the marketing director believes that the current volume of 15 million units could be maintained despite an increase in price of 50¢ per unit. On this basis, calculate the implied arc price elasticity of demand. C. Holding all else equal, would a further increase in price result in higher or lower total revenue?
1
Expert's answer
2017-10-13T14:57:06-0400
a) EI = ((Q1-Q)/Q)/((I1/I)/I) = ((25-15)/15)/((18,75-17,25)/17,25)=7,7.
b) Ep = ((Q1-Q)/Q)/((P1/P)/P) = ((15-15)/15)/((7,75-8,25)/7,75)=0.
c) I believe that an increase in the price of this product will lead to an increase in revenues because the economy is coming out of the crisis, so the amount of purchased products will not change especially with the increase in prices.